# High-dimensional Autoregressive Modeling for Time Series with Hierarchical Structures

### v1

@@ 3.2 Non-asymptotic properties for tensor regression
We begin with the tensor-on-tensor regression setting, where $\{(\mbox{$\mathscr{Y}$}_{t},\mbox{$\mathscr{X}$}_{t})\}_{t=1}^{T}$ are assumed to be independent and identically distributed (i.i.d.) observations. The following assumptions are imposed.
@@
Predictors $\{\boldsymbol{x}_{t}\}$ are i.i.d. with $\mathbb{E}(\boldsymbol{x}_{t})=\boldsymbol{0}$, $\mathrm{var}(\boldsymbol{x}_{t})=\boldsymbol{\Sigma}_{x}$, and follow a $\sigma^{2}$-sub-Gaussian distribution. There exist constants $0<c_{x}<C_{x}<\infty$ such that $c_{x}\leq\lambda_{\min}(\boldsymbol{\Sigma}_{x})\leq\lambda_{\max}(\boldsymbol{\Sigma}_{x})\leq C_{x}$, where the two quantities $c_{x}$ and $C_{x}$ depend on the dimensions of $p_{j}$’s, and they may shrink to zero or diverge to infinity as the dimensions increase.
@@
Error terms $\{\boldsymbol{e}_{t}\}$ are i.i.d. and, conditional on $\boldsymbol{x}_{t}$, $\boldsymbol{e}_{t}$ has mean zero and follows $\kappa^{2}$-sub-Gaussian distribution.
@@
The matrix $\boldsymbol{\Theta}_{\mathrm{LR}}$ has its operator norm upper bounded by a constant $g_{1}>0$.
Assumption 2 guarantees the well-conditioning of the covariance matrix of the predictors. Assumption 3, on the other hand, controls the tail behavior of the noise by assuming sub-Gaussianity, a standard requirement in high-dimensional statistics that facilitates concentration inequalities. Both assumptions are widely adopted in the literature on high-dimensional regression and low-rank estimation; see, for example, wainwright2019high; si2024efficient; cai2025efficient. Assumption 4 imposes an upper bound on the operator norm of the coefficient matrix $\boldsymbol{\Theta}_{\mathrm{LR}}$. This is a mild technical condition, as the dimensions $s_{\mathrm{tot}}$ and $r_{\mathrm{tot}}$ of $\boldsymbol{\Theta}_{\mathrm{LR}}$ are typically small by carefully choosing the action orders.
The following theorem provides a non-asymptotic upper bound on the estimation error of estimator $\widehat{\mbox{$\mathscr{A}$}}_{\mathrm{LR}}$.
@@
Suppose that Assumptions 2 - 4 hold, and the true coefficient tensor $\mbox{$\mathscr{A}$}_{\mathrm{LR}}^{*}\in\mathcal{S}_{\mathrm{LR}}$. If the sample size $T\gtrsim(\sigma^{4}/c_{x}^{2})d_{\mathrm{LR}}\log((\sigma^{4}/c_{x}^{2})d_{\mathrm{LR}})$, then there exists $c_{1}>0$ such that
@@
with probability at least $1-2\exp\left\{-c_{1}d_{\mathrm{LR}}\log T\right\}$, where $d_{\mathrm{LR}}=\sum_{k=1}^{K_{x}}d_{x}(\alpha_{k})+\sum_{k=1}^{K_{y}}d_{y}(\beta_{k})+sr$.
We define $d_{\mathrm{LR}}$ as the model complexity of the low-rank regression estimator, which can be rewritten explicitly as $d_{\mathrm{LR}}=K_{x}r_{\max}^{2}\sum_{m=1}^{M}p_{m}+K_{y}s_{\max}^{2}\sum_{n=1}^{N}q_{n}+K_{x}K_{y}r_{\max}s_{\max}$, where $r_{\max}\vcentcolon=\max_{m,k}r_{m,k}$ and $s_{\max}\vcentcolon=\max_{n,k^{\prime}}s_{n,k^{\prime}}$. From Theorem 1, our proposed low-rank regression estimator $\widehat{\mbox{$\mathscr{A}$}}_{\mathrm{LR}}$, up to a logarithmic factor, achieves an estimation error bound of order $O_{p}(\sqrt{d_{\mathrm{LR}}/T})$ when the sample size $T\gtrsim d_{\mathrm{LR}}/\text{SNR}^{2}$, where $\text{SNR}=c_{x}/\sigma^{2}$ is the predictor signal-to-noise ratio. This rate aligns with the general optimal rate for low-rank regression established in han2022optimal. The model complexity $d_{\mathrm{LR}}$ grows linearly with the dimensions of each mode $\{p_{m}\}$ and $\{q_{n}\}$, and quadratically with the interim ranks $\{r_{m,k}\}$ and $\{s_{n,k^{\prime}}\}$. The logarithmic factor $\log T$, negligible in most common cases, arises from the non-orthogonality of the loadings. In contrast, if we assume the coefficient tensor $\mathscr{A}$ follows a Tucker decomposition with Tucker ranks $(s_{1},s_{2},\ldots,s_{N},r_{1},r_{2},\ldots,r_{M})$, the corresponding complexity would be $d_{\mathrm{Tucker}}=\prod_{n=1}^{N}s_{n}\times\prod_{m=1}^{M}r_{m}+\sum_{n=1}^{N}q_{n}s_{n}+\sum_{m=1}^{M}p_{m}r_{m}$, which grows exponentially with the tensor orders $N$ and $M$ (han2022optimal; wang2022high).
3.3 Non-asymptotic properties for tensor autoregression
We now turn to the tensor autoregressive model specified in (2.12). This setting introduces temporal dependence, and the assumptions are adapted accordingly.
@@ 3.3 Non-asymptotic properties for tensor autoregression
Let error terms $\boldsymbol{e}_{t}=\boldsymbol{\Sigma}_{e}^{1/2}\boldsymbol{\xi}_{t}$. Random vectors $\{\boldsymbol{\xi}_{t}\}$ are i.i.d. with mean zero and covariance $\boldsymbol{I}_{Q}$, the entries $(\boldsymbol{\xi}_{tj})_{1\leq j\leq Q}$ of $\boldsymbol{\xi}_{t}$ are mutually independent and $\kappa^{2}$-sub-Gaussian distributed, and there exist constants $0<c_{e}<C_{e}<\infty$ such that $c_{e}\leq\lambda_{\min}(\boldsymbol{\Sigma}_{e})\leq\lambda_{\max}(\boldsymbol{\Sigma}_{e})\leq C_{e}$.
@@
The matrix $\boldsymbol{\Theta}_{\mathrm{AR}}$ has its operator norm upper bounded by a constant $g_{2}>0$.
The sub-Gaussianity assumption in Assumption 5 is commonly used in the analysis of high-dimensional time series models. It provides control over the tail behavior of the error terms and facilitates the use of concentration inequalities. This assumption is widely adopted in recent literature, including zheng2020finite; wang2024high; huang2025supervised. Assumption 6 is essential for the consistency of the estimator in the autoregressive setting, which is relatively mild and practically reasonable. In our setting, $\boldsymbol{\Lambda}_{y}$ and $\boldsymbol{\Lambda}_{x}$ may exhibit multi-collinearity, and this assumption is weaker than enforcing a lower bound on the smallest singular value of $\boldsymbol{\Lambda}_{x}$ and $\boldsymbol{\Lambda}_{y}$ for model well-behavedness.
Following basu2015regularized, we define the following quantities,
@@
where $\bar{\mbox{$\mathscr{A}$}}(z)$ denotes the conjugate transpose of $\mbox{$\mathscr{A}$}(z)$.
Furthermore, the AR($L$) model in (2.12) can be alternatively represented as a $QL$-dimensional AR(1) model with its transition matrix $\boldsymbol{B}$ defined in (C.3) in the Supplementary Material. We similarly define the matrix polynomial associated with the lifted AR(1) representation as $\boldsymbol{B}(z)=\boldsymbol{I}_{QL}-\boldsymbol{B}z$. We then define
$\mu_{\mathrm{min}}(\boldsymbol{B})=\min_{|z|=1}\lambda_{\mathrm{min}}(\bar{\boldsymbol{B}}(z)\boldsymbol{B}(z))\quad\text{and}\quad\mu_{\mathrm{max}}(\boldsymbol{B})=\max_{|z|=1}\lambda_{\mathrm{max}}(\bar{\boldsymbol{B}}(z)\boldsymbol{B}(z)),$
which quantify the smallest and largest eigenvalues the Hermitian matrix $\bar{\boldsymbol{B}}(z)\boldsymbol{B}(z)$ can take on the unit circle. Note that $\mu_{\mathrm{min}}(\boldsymbol{B})$, $\mu_{\mathrm{max}}(\boldsymbol{B})$ are not necessarily the same as $\mu_{\mathrm{min}}(\mbox{$\mathscr{A}$})$, $\mu_{\mathrm{max}}(\mbox{$\mathscr{A}$})$ (basu2015regularized), and we can show $\mu_{\mathrm{min}}(\boldsymbol{B}^{\ast})\leq\mu_{\mathrm{max}}(\mbox{$\mathscr{A}$}^{\ast})$.
We then state the error upper bound for the estimator $\widehat{\mbox{$\mathscr{A}$}}_{\mathrm{AR}}$ below.
@@
Suppose that Assumptions 1, 5, and 6 hold, and the true coefficient tensor $\mbox{$\mathscr{A}$}_{\mathrm{AR}}^{\ast}\in\mathcal{S}_{\mathrm{AR}}$. If the sample size $T\gtrsim(\kappa^{4}\kappa_{U,B}^{2}/\kappa_{L,A}^{2})d_{\mathrm{AR}}\log((\kappa^{4}\kappa_{U,B}^{2}/\kappa_{L,A}^{2})d_{\mathrm{AR}})$, then
@@
with probability at least $1-2\exp\left\{-c_{1}d_{\mathrm{AR}}\log T\right\}$, where $d_{\mathrm{AR}}=\sum_{k=1}^{K_{x}}d_{x}(\alpha_{k})+\sum_{k=1}^{K_{y}}d_{y}(\beta_{k})+Lsr$, $\boldsymbol{B}^{\ast}$ is the true value of $\boldsymbol{B}$ constructed from $\mbox{$\mathscr{A}$}^{\ast}$, $\kappa_{U,B}=C_{e}/\mu_{\mathrm{min}}(\boldsymbol{B}^{\ast})$, $\kappa_{L,A}=c_{e}/\mu_{\mathrm{max}}(\mbox{$\mathscr{A}$}^{\ast})$, $\kappa^{\prime}_{U,B}=C_{e}/\mu^{1/2}_{\mathrm{min}}(\boldsymbol{B}^{\ast})$, and $c_{1}>0$ is a constant.
We define $d_{\mathrm{AR}}$ as the model complexity of the low-rank autoregressive estimator, which can be explicitly expressed as $d_{\mathrm{AR}}=(K_{x}r_{\max}^{2}+K_{y}s_{\max}^{2})\sum_{n=1}^{N}q_{n}+LK_{x}K_{y}r_{\max}s_{\max}$. When the order $N$, the number of action orders $K_{x}$, $K_{y}$, and the largest ranks $s_{\max}$ and $r_{\max}$ are held fixed, the key parameters $c_{e}$, $C_{e}$, $\kappa^{2}$, $\mu_{\mathrm{min}}(\boldsymbol{B}^{\ast})$, and $\mu_{\mathrm{max}}(\mbox{$\mathscr{A}$}^{\ast})$ are all bounded away from zero and infinity. Consequently, the estimation error of the low-rank tensor autoregressive estimator satisfies $\|{\mbox{$\mathscr{\widehat{A}}$}_{\mathrm{AR}}-\mbox{$\mathscr{A}$}_{\mathrm{AR}}^{\ast}}\|_{\mathrm{F}}=O_{p}(\sqrt{d_{\mathrm{AR}}/T})$ up to a logarithmic term, where $d_{\mathrm{AR}}$ is the effective model complexity of the autoregressive estimator.
@@ 4.1 Algorithms
This subsection introduces an algorithm for computing the low-rank estimator defined in (3.1) under the general setting. To preserve the sequential low-rank structure of $\mbox{$\mathscr{A}$}]_{N}$, we adopt an alternating least squares (ALS) approach, which iteratively updates each component $\boldsymbol{H}_{n,k^{\prime}}$ and $\boldsymbol{G}_{m,k}$ while keeping all other components fixed.
We introduce some shorthand notations for clarity. Let $\boldsymbol{y}_{t}=\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})$ and $\boldsymbol{x}_{t}=\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})=(\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t-1})^{\top},\dots,\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t-L})^{\top})^{\top}$, and our model can be rewritten in the vectorized form as:
$\boldsymbol{y}_{t}=\boldsymbol{\Lambda}_{y}\boldsymbol{\Theta}(\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x}^{\top})\boldsymbol{x}_{t}+\boldsymbol{e}_{t},$ (4.1)
@@
We sequentially update each orthogonal matrix in each row block of $\boldsymbol{\Lambda}_{x}^{\top}$ (i.e., update $\boldsymbol{G}_{1,1}$ up to $\boldsymbol{G}_{M,1}$ for the first row block), and then switch to the next row block for another sequence of updates, and so on. Specifically, we convert the update of each component into a regression problem relegated to Supplementary Material, similar to coordinate descent. The update for $\boldsymbol{\Lambda}_{y}^{\top}$ follows a similar procedure. After updating all components in $\boldsymbol{\Lambda}_{x}^{\top}$ and $\boldsymbol{\Lambda}_{y}^{\top}$, we orthogonalize each row block of the two loading matrices and update the core tensor $\boldsymbol{\Theta}$ via an SSVD algorithm detailed in Supplementary Material. This process is repeated until convergence. For convenience, we arrange these matrices into $\boldsymbol{L}_{1},\boldsymbol{L}_{2},\ldots,\boldsymbol{L}_{K_{x}M+K_{y}N}$, where $\boldsymbol{L}_{i}$ is the $i$-th orthogonal matrix in the update procedure, and we re-define the loss function with regard to $\boldsymbol{L}_{i}$’s and $\boldsymbol{\Theta}$ as follows:
$\mathcal{L}_{T}\big(\boldsymbol{L}_{1},\dots,\boldsymbol{L}_{K_{x}M+K_{y}N},\boldsymbol{\Theta}\big)=\frac{1}{T}\sum_{t=1}^{T}\|\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})-\boldsymbol{\Lambda}_{y}\boldsymbol{\Theta}(\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x}^{\top})\operatorname*{vec}{(\mbox{$\mathscr{X}$}_{t})}\|^{2},$
$\text{s.t., }\boldsymbol{G}_{m,k}=\boldsymbol{L}_{(k-1)M+m}\quad\text{and}\quad\boldsymbol{H}_{n,k^{\prime}}=\boldsymbol{L}_{K_{x}M+(k^{\prime}-1)N+n}$ (4.2)
for all $m,k,n,k^{\prime}$, and $\boldsymbol{\Lambda}_{x}$ and $\boldsymbol{\Lambda}_{y}$ are defined in (2.10), respectively. The overall algorithm is summarized below:
Algorithm 1 ALS for low-rank estimation
Input: Data $\{(\boldsymbol{y}_{t},\boldsymbol{x}_{t})\}_{t=1}^{T}$, initial coefficient $\mbox{$\mathscr{A}$}^{(0)}]_{N}=\boldsymbol{\Lambda}_{y}^{(0)}\boldsymbol{\Theta}^{(0)}\boldsymbol{\Lambda}_{x}^{(0)\top}$, and components $\{\{\boldsymbol{H}^{(0)}_{n,k^{\prime}}\}_{n=1}^{N}\}_{k^{\prime}=1}^{K_{y}}$ and $\{\{\boldsymbol{G}^{(0)}_{m,k}\}_{m=1}^{M}\}_{k=1}^{K_{x}}$ arranged into $\{\boldsymbol{L}_{i}^{(0)}\}_{i=1}^{K_{x}M+K_{y}N}$.
repeat
  for $i=1,2,\dots,K_{x}M+K_{y}N$ do
   $\boldsymbol{L}^{(u+0.5)}_{i}=\operatorname*{arg\,min}_{\boldsymbol{L}_{i}}\mathcal{L}_{T}\big(\boldsymbol{L}_{1}^{(u+0.5)},\dots,\boldsymbol{L}_{i-1}^{(u+0.5)},\boldsymbol{L}_{i},\boldsymbol{L}_{i+1}^{(u)},\dots,\boldsymbol{L}_{K_{x}M+K_{y}N}^{(u)},\boldsymbol{\Theta}^{(u)}\big)$
  end for
  $\{\boldsymbol{L}_{i}^{(u+1)}\}_{i=1}^{K_{x}M+K_{y}N},\boldsymbol{\Theta}^{(u)}=\text{SSVD}(\{\boldsymbol{L}_{i}^{(u+0.5)}\}_{i=1}^{K_{x}M+K_{y}N},\boldsymbol{\Theta}^{(u)})$
  $\boldsymbol{\Theta}^{(u+1)}=\operatorname*{arg\,min}_{\boldsymbol{\Theta}}\mathcal{L}_{T}\big(\boldsymbol{L}_{1}^{(u+1)},\dots,\boldsymbol{L}_{K_{x}M+K_{y}N}^{(u+1)},\boldsymbol{\Theta}\big)$
until convergence or $u=u_{\max}$, where $u_{\max}$ is the maximum iteration
Return $\mbox{$\mathscr{A}$}^{(u)}]_{N}=\boldsymbol{\Lambda}_{y}^{(u)}\boldsymbol{\Theta}^{(u)}\boldsymbol{\Lambda}_{x}^{(u)\top}$ via (4.2), (2.10).

### v2

@@ 3.2 Non-asymptotic properties for tensor regression settings
Predictor $\boldsymbol{x}_{t}$ follows a $\sigma^{2}$-sub-Gaussian distribution with $\mathbb{E}(\boldsymbol{x}_{t})=\boldsymbol{0}$ and $\mathrm{var}(\boldsymbol{x}_{t})=\boldsymbol{\Sigma}_{x}$. There exist constants $0<c_{x}<C_{x}<\infty$ such that $c_{x}\leq\lambda_{\min}(\boldsymbol{\Sigma}_{x})\leq\lambda_{\max}(\boldsymbol{\Sigma}_{x})\leq C_{x}$, where the two quantities $c_{x}$ and $C_{x}$ depend on the dimensions of $p_{j}$’s, and they may shrink to zero or diverge to infinity as the dimensions increase.
@@
Error term $\boldsymbol{e}_{t}$, conditional on predictor $\boldsymbol{x}_{t}$, follows $\kappa^{2}$-sub-Gaussian distribution with mean zero.
@@
It holds that $\|\boldsymbol{\Theta}_{\mathrm{LR}}\|_{\mathrm{op}}<g_{1}$, where $\|\cdot\|_{\mathrm{op}}$ is the operator norm.
Assumption 2 guarantees the well-conditioning of predictors, while Assumption 3 controls the tail behavior of noises by the sub-Gaussianity. The two assumptions are standard high-dimensional settings, and they have been widely used in the literature (Wainwright, 2019; Si et al., 2024; Cai et al., 2025). Assumption 4 imposes an upper bound on the operator norm of the coefficient matrix $\boldsymbol{\Theta}_{\mathrm{LR}}\in\mathbb{R}^{s\times r}$. It is used to handle the non-orthogonality of loading matrices $\boldsymbol{\Lambda}_{x}$ and $\boldsymbol{\Lambda}_{y}$, and we can remove it with the cost of a large model complexity. This assumption is mild since both dimensions $s$ and $r$ are typically small by carefully choosing the action orders. Denote the model complexity by $d_{\mathrm{LR}}=\sum_{k=1}^{K_{x}}d_{x}(\alpha_{k})+\sum_{k=1}^{K_{y}}d_{y}(\beta_{k})+sr$, and the estimation error bound is stated below.
@@
Suppose that Assumptions 2 - 4 hold, and the true coefficient tensor $\mbox{$\mathscr{A}$}_{\mathrm{LR}}^{*}\in\mathcal{S}_{\mathrm{LR}}$. If the sample size $T\gtrsim(\sigma^{4}/c_{x}^{2})d_{\mathrm{LR}}\log\{(\sigma^{4}/c_{x}^{2})d_{\mathrm{LR}}\}$, then,
@@
with probability at least $1-2\exp\left\{-C_{1}d_{\mathrm{LR}}\log T\right\}$, where $C_{1}$ is a positive constant.
Let $r_{\max}\vcentcolon=\max_{m,k}r_{m,k}$ and $s_{\max}\vcentcolon=\max_{n,k^{\prime}}s_{n,k^{\prime}}$, and the model complexity can be simplified as $d_{\mathrm{LR}}\leq K_{x}r_{\max}^{2}\sum_{m=1}^{M}p_{m}+K_{y}s_{\max}^{2}\sum_{n=1}^{N}q_{n}+K_{x}K_{y}r_{\max}s_{\max}$. From the above theorem, the OLS estimator $\widehat{\mbox{$\mathscr{A}$}}_{\mathrm{LR}}$ achieves an estimation error bound of order $O_{p}(\sqrt{d_{\mathrm{LR}}/T})$, up to a logarithmic factor, when the sample size $T\gtrsim d_{\mathrm{LR}}/(c_{x}/\sigma^{2})^{2}$. This rate aligns with the general optimal rate for low-rank regression established in Han et al. (2022). The model complexity $d_{\mathrm{LR}}$ grows linearly with the dimensions of each mode $\{p_{m}\}$ and $\{q_{n}\}$, and quadratically with the interim ranks $\{r_{m,k}\}$ and $\{s_{n,k^{\prime}}\}$.
3.3 Non-asymptotic properties for tensor autoregressive models
This subsection states non-asymptotic properties of the OLS estimator $\mbox{$\mathscr{\widehat{A}}$}_{\mathrm{AR}}$ for tensor autoregressive models at (2.12) and (2.13), and we first introduce some conditions below.
@@ 3.3 Non-asymptotic properties for tensor autoregressive models
Let $\boldsymbol{e}_{t}=\boldsymbol{\Sigma}_{e}^{1/2}\boldsymbol{\xi}_{t}$ with $\boldsymbol{\Sigma}_{e}=\mathrm{var}(\boldsymbol{e}_{t})$. Random vectors $\{\boldsymbol{\xi}_{t}\}$ are $i.i.d.$ with mean zero and covariance matrix $\boldsymbol{I}_{Q}$, and the entries of $\boldsymbol{\xi}_{t}\in\mathbb{R}^{Q}$ are independent and $\kappa^{2}$-sub-Gaussian distributed. Moreover, there exist constants $0<c_{e}<C_{e}<\infty$ such that $c_{e}\leq\lambda_{\min}(\boldsymbol{\Sigma}_{e})\leq\lambda_{\max}(\boldsymbol{\Sigma}_{e})\leq C_{e}$, where the two quantities depend on the dimensions of $q_{j}$’s, and they may shrink to zero or diverge to infinity as the dimensions increase.
@@
It holds that $\|\boldsymbol{\Theta}_{\mathrm{AR}}\|_{\mathrm{op}}<g_{2}$, where $\|\cdot\|_{\mathrm{op}}$ is the operator norm.
The sub-Gaussian assumption has also been commonly used in the literature of high-dimensional time series (Zheng and Cheng, 2020; Wang et al., 2024; Huang et al., 2025).
We next establish non-asymptotic properties of the OLS estimator $\mbox{$\mathscr{\widehat{A}}$}_{\mathrm{AR}}$, and they rely on the temporal and cross-sectional dependence of $\{\mbox{$\mathscr{Y}$}_{t}\}$ (Basu and Michailidis, 2015). To this end, for the matrix polynomial of model (2.12), we define the quantities below,
@@
[…]
@@
Suppose that Assumptions 1, 5 and 6 hold, and the true coefficient tensor $\mbox{$\mathscr{A}$}_{\mathrm{AR}}^{\ast}\in\mathcal{S}_{\mathrm{AR}}$. If the sample size $T\gtrsim(\kappa^{4}\kappa_{U,B}^{2}/\kappa_{L,A}^{2})d_{\mathrm{AR}}\log\{\kappa^{4}\kappa_{U,B}^{2}/\kappa_{L,A}^{2})d_{\mathrm{AR}}\}$, then
@@
with probability at least $1-2\exp\left\{-C_{1}d_{\mathrm{AR}}\log T\right\}$, where $C_{1}>0$ is a positive constant.
[…]
@@ 4.1 Algorithms
This subsection introduces an alternating least squares (ALS) method to search for the estimate $\widehat{\mbox{$\mathscr{A}$}}_{\mathrm{AR}}$ in Section 3.1. Specifically, for each $\mbox{$\mathscr{A}$}\in\mathcal{S}_{\mathrm{AR}}$ at (3.1), its mode-$N$ sequential matricization has the form of $\mbox{$\mathscr{A}$}]_{N}=\boldsymbol{\Lambda}_{y}\boldsymbol{\Theta}(\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x}^{\top})$, where $\boldsymbol{\Theta}=(\boldsymbol{\Theta}_{1},\ldots,\boldsymbol{\Theta}_{L})$. Let
$\boldsymbol{\Lambda}_{x,k}$ $=\boldsymbol{T}_{x}^{\top}(\alpha_{k})(\boldsymbol{I}_{p_{\alpha^{k}_{(2)}}\cdots p_{\alpha^{k}_{(M)}}}\otimes\boldsymbol{G}_{1,k})\cdots(\boldsymbol{I}_{p_{\alpha^{k}_{(M)}}}\otimes\boldsymbol{G}_{M-1,k})\boldsymbol{G}_{M,k}\in\mathbb{R}^{P\times r_{M,k}}\hskip 8.53581pt\text{and}$
$\boldsymbol{\Lambda}_{y,k^{\prime}}$ $=\boldsymbol{T}_{y}^{\top}(\beta_{k^{\prime}})(\boldsymbol{I}_{q_{\beta^{k^{\prime}}_{(2)}}\cdots q_{\beta^{k^{\prime}}_{(N)}}}\otimes\boldsymbol{H}_{1,k^{\prime}})\cdots(\boldsymbol{I}_{q_{\beta^{k^{\prime}}_{(N)}}}\otimes\boldsymbol{H}_{N-1,k^{\prime}})\boldsymbol{H}_{N,k^{\prime}}\in\mathbb{R}^{Q\times s_{N,k^{\prime}}}$
for $1\leq k\leq K_{x}$ and $1\leq k^{\prime}\leq K_{y}$, and it then holds that $\boldsymbol{\Lambda}_{x}=(\boldsymbol{\Lambda}_{x,1},\ldots,\boldsymbol{\Lambda}_{x,K_{x}})\in\mathbb{R}^{P\times r}$ and $\boldsymbol{\Lambda}_{y}=(\boldsymbol{\Lambda}_{y,1},\ldots,\boldsymbol{\Lambda}_{y,K_{y}})\in\mathbb{R}^{Q\times s}$; see (2.10) for details. Moreover, for each $1\leq l\leq L$, $\mbox{$\mathscr{A}$}_{l}]_{N}=\boldsymbol{\Lambda}_{y}\boldsymbol{\Theta}_{l}\boldsymbol{\Lambda}_{x}^{\top}=\sum_{k^{\prime}=1}^{K_{y}}\sum_{k=1}^{K_{x}}\boldsymbol{\Lambda}_{y,k^{\prime}}\boldsymbol{\Theta}_{k^{\prime},k,l}\boldsymbol{\Lambda}_{x,k}^{\top}$, where $\boldsymbol{\Theta}_{k^{\prime},k,l}\in\mathbb{R}^{s_{N,k^{\prime}}\times r_{M,k}}$ with $1\leq k\leq K_{x}$ and $1\leq k^{\prime}\leq K_{y}$ are the corresponding submatrices of $\boldsymbol{\Theta}_{l}$. As a result, it is equivalent to minimizing the loss function of $\mathcal{L}_{T}(\mbox{$\mathscr{A}$})$ at (3.1) with respect to two blocks of parameters: $\{\boldsymbol{\Lambda}_{x,k}(\boldsymbol{G}_{m,k},1\leq m\leq M),1\leq k\leq K_{x}\}$, $\{\boldsymbol{\Lambda}_{y,k}(\boldsymbol{H}_{n,k},1\leq n\leq N),1\leq k^{\prime}\leq K_{y}\}$, and $\boldsymbol{\Theta}$, and the first two blocks of parameters correspond to $\boldsymbol{\Lambda}_{x}$ and $\boldsymbol{\Lambda}_{y}$, respectively.
Algorithm 1 ALS for tensor autoregression
Input: Time series $\{\boldsymbol{y}_{t},-L+1\leq t\leq T\}$, initial coefficients $\mbox{$\mathscr{A}$}^{(0)}]_{N}=\boldsymbol{\Lambda}_{y}^{(0)}\boldsymbol{\Theta}^{(0)}(\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x}^{(0)\top})$ with component matrices $\{\{\boldsymbol{H}^{(0)}_{n,k}\}_{n=1}^{N}\}_{k=1}^{K_{y}}$ and $\{\{\boldsymbol{G}^{(0)}_{m,k}\}_{m=1}^{M}\}_{k=1}^{K_{x}}$.
repeat
  for $k=1,2,\ldots,K_{x}$ do
   Sequentially update $\boldsymbol{G}_{m,k}$ for $1\leq m\leq M$ while keeping all other components being fixed at the values from the last iteration.
   Standardize each $\boldsymbol{G}_{m,k}$ for $1\leq m\leq M$ such that each column has unit norm.
  end for
  for $k=1,2,\ldots,K_{y}$ do
   Sequentially update $\boldsymbol{H}_{n,k}$ for $1\leq n\leq N$ while keeping all other components being fixed at the values from the last iteration.
   Standardize each $\boldsymbol{H}_{n,k}$ for $1\leq n\leq N$ such that each column has unit norm.
  end for
  Update $\boldsymbol{\Theta}=(\boldsymbol{\Theta}_{1},\ldots,\boldsymbol{\Theta}_{L})$.
until convergence
 Call the SSVD algorithm to orthogonalize all $\boldsymbol{G}_{m,k}$’s and $\boldsymbol{H}_{n,k}$’s, and then update $\boldsymbol{\Theta}$.
Return $\widehat{\mbox{$\mathscr{A}$}}_{\mathrm{AR}}$ such that $\widehat{\mbox{$\mathscr{A}$}}_{\mathrm{AR}}]_{N}=\boldsymbol{\Lambda}_{y}\boldsymbol{\Theta}(\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x}^{\top})$.
Let $\boldsymbol{y}_{t}=\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})$, and the tensor autoregressive model at (2.12) and (2.13) has the form of $\boldsymbol{y}_{t}=\boldsymbol{\Lambda}_{y}\boldsymbol{\Theta}_{1}\boldsymbol{\Lambda}_{x}^{\top}\boldsymbol{y}_{t-1}+\cdots+\boldsymbol{\Lambda}_{y}\boldsymbol{\Theta}_{L}\boldsymbol{\Lambda}_{x}^{\top}\boldsymbol{y}_{t-L}+\boldsymbol{e}_{t}$ or
$\begin{split}\boldsymbol{y}_{t}=\sum_{k=1}^{K_{x}}\sum_{k^{\prime}=1}^{K_{y}}\boldsymbol{\Lambda}_{y,k^{\prime}}\boldsymbol{\Theta}_{k,k^{\prime},1}\boldsymbol{\Lambda}_{x,k}^{\top}\boldsymbol{y}_{t-1}+\cdots+\sum_{k=1}^{K_{x}}\sum_{k^{\prime}=1}^{K_{y}}\boldsymbol{\Lambda}_{y,k^{\prime}}\boldsymbol{\Theta}_{k,k^{\prime},L}\boldsymbol{\Lambda}_{x,k}^{\top}\boldsymbol{y}_{t-L}+\boldsymbol{e}_{t},\end{split}$ (4.1)
@@
In addition, when updating $\boldsymbol{\Lambda}_{x,k}(\boldsymbol{G}_{m,k},1\leq m\leq M)$ for each $1\leq k\leq K_{x}$, we sequentially update each of its component matrices $\boldsymbol{G}_{m,k}$ with $1\leq m\leq M$. Interestingly, the update of each $\boldsymbol{G}_{m,k}$ can be formulated as a least squares problem by fixing all other components, and it can be verified to have a solution with closed form; see Section B.2 of the supplementary file for details. The situation is the same when sequentially updating the component matrices of $\boldsymbol{\Lambda}_{y,k}(\boldsymbol{H}_{n,k},1\leq n\leq N)$ for each $1\leq k\leq K_{y}$, and hence all updating steps at Algorithm 1 have solutions with closed forms. Moreover, these component matrices are all orthogonal with unit norm, i.e., they are orthonormal. To guarantee the convergence, we give up the orthogonality during iterations, and these component matrices will be orthogonalized at the end by a SSVD algorithm detailed at Section B.4 of the supplementary file.
For the initialization, we may simply consider random numbers around zero. In addition, the algorithm to search for the estimate $\widehat{\mbox{$\mathscr{A}$}}_{\mathrm{LR}}$ in Section 3.1 can be designed similarly.